You are making an electric circuit with a battery providing 3.0 V, and two resistors of resistance of 2.0 Ohm and 8.0 Ohm

connected in parallel to the battery. What is the current out of the battery in A? 

knowing the potential difference across both resistors is 3.0 V, calculate the current going through the

2.0-Ohm resistor in A.


1
Expert's answer
2021-12-15T11:55:19-0500

The current out of the battery


"R=\frac{R_1R_2}{R_1+R_2}=\frac{2\cdot 8}{2+8}=1.6\ (\Omega)"


"I=V/R=3/1.6=1.875\ (A)"


The current going through the "2.0\ Ohm" resistor


"I_1=V/R_1=3/2=1.5\ (A)"